### v1

# When Large Language Models Meet UAVs: How Far Are We?

Abstract.
The integration of unmanned aerial vehicles (UAVs) and large language models (LLMs) has emerged as a research direction of growing interest, with the potential to address challenges in autonomous decision-making, human-UAV interaction, and real-time adaptability. However, existing studies have remained largely in preliminary exploration with a limited understanding of real-world practice, risking a misalignment between academic research and practical needs and hindering the translation of results. To examine and address these potential challenges, we conducted an empirical study of 74 selected papers and 56 public GitHub projects, identified nine task types for LLMs in UAV systems, and quantified their distribution. Our findings show that academic research emphasizes theoretical modeling and task optimization with dispersed attention across tasks. In contrast, industrial projects focus on flight control, task planning, and human-machine interaction, prioritizing operability and efficiency. To further capture industry perspectives, we distributed an online questionnaire. We obtained 52 valid responses: 40.4% of practitioners have attempted to apply LLMs to UAV tasks. We further identify factors that impede real-world integration, including technological maturity, performance, safety, cost, and other considerations. Finally, we highlight challenges for future development and provide recommendations.
Large Language Models, Unmanned Aerial Vehicles, Empirical Study
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software system structuresccs: Software and its engineering Designing softwareccs: Computer systems organization Embedded and cyber-physical systems
[…]
1. Introduction
In recent years, unmanned aerial vehicles (UAVs) have been rapidly applied in agriculture, disaster response, security inspection, logistics, and transportation, offering significant efficiency gains due to their mobility, low deployment cost, and diverse capabilities (Shakhatreh et al., 2019; Liu et al., 2021; Colomina and Molina, 2014). However, while UAV operations involve complex processes of environmental perception, flight control, and task planning, current UAV systems still primarily rely on manual operation or rule-based pipelines (Gupta et al., 2016), which limits their flexibility and adaptability in dynamic environments and diverse tasks. Consequently, it is still difficult to achieve scalable high-autonomy operations and thereby constraining the broader deployment of UAVs in more complex scenarios (Colomina and Molina, 2014).
Advances in large language models (LLMs) in language understanding, reasoning, and generation (Chang et al., 2024; Zhao et al., 2023a) offer new possibilities to address the above limitations. By enabling UAVs to incorporate reasoning and interpretation in complex environments, LLMs reduce reliance on manual operation and rule-based pipelines, allowing tasks to be decomposed and decisions to be made with greater flexibility (Tian et al., 2025; Wang et al., 2024a; Zeng et al., 2023a; Ahn et al., 2022; Brohan et al., 2023). This capability can significantly enhance the autonomy and adaptability of UAVs, making the integration of UAVs with LLMs an increasingly prominent research direction.
Although existing studies (Javaid et al., 2024; Tian et al., 2025; TAZIR et al., 2023) demonstrate the feasibility of this direction, most remain confined to simulation or controlled environments, with limited coverage of scenarios and task diversity. They often give insufficient consideration to critical factors in real-world applications, such as computational resources, latency, reliability, and compliance (Zhang et al., 2023; Wu et al., 2024), and the defined task metrics do not readily correspond to the complexity of practical demands. These limitations risk creating a mismatch between academic research and actual needs, thereby reducing the practical value of research studies in translating them into technologies usable by industry.
To bridge the gap, this paper systematically explores the following three key research questions (RQs):
RQ1. What is the current research status and focus of large language models in the field of UAVs?
Identifying the specific tasks that LLMs can perform in UAV systems is the first step toward understanding the advantages and limitations of their integration. To this end, we conducted a systematic literature review (SLR) (Kitchenham et al., 2009) and analyzed a total of 997 research papers related to LLMs and UAVs, applying a card sorting method (Spencer, 2009) to summarize and classify their contents. Based on the SLR process, we classified the UAV-LLM integration into four core stages in the UAV workflow—Information Input, Task Planning, Control Execution, and Result Feedback—each corresponding to a set of LLM-enabled tasks. Among these research directions, the top three most frequently studied directions are Single-UAV Task Reasoning and Planning, Control Logic and Command Generation, and Natural Language Command Parsing. Our characterization of UAV-LLM research clarifies how LLMs are currently applied within UAV systems, which also enables us to compare academic research with industrial practices in subsequent analysis.
RQ2. What are the differences between academia and industry in integrating UAVs with large language models?
Due to differences in objectives, resources, and evaluation standards, academia and industry often emphasize different aspects when using new technologies. In this study, we aim to compare these differences and explore how the two sides can be better connected, so that research results are more practical and industrial needs can be more clearly reflected in future research. We collected 1,509 UAV-LLM-related projects from GitHub and performed an empirical analysis of their characteristics and development practices. Our analysis revealed different efforts between academia and industry in the UAV-LLM field. Academic research focuses more on the design of complex system integration and theoretical mechanisms. It is committed to exploring how to use LLMs to improve system autonomy, flexibility, and multi-tasking capabilities. In contrast, industrial practice emphasizes system stability, low-cost implementation, and rapid deployment, and generally tends to adopt efficient solutions. This, to some extent, limits the in-depth application of LLMs in scenarios requiring high timeliness or operating under resource constraints. By revealing the gap between academia and industry, our results could provide practical implications for facilitating real-world UAV-LLM practices.
RQ3. How do real-world practitioners perceive these differences, and what factors contribute to them?
Although our comparative analysis reveals the task preferences of UAV-LLM in different scenarios, the root causes of these differences remain unclear. To this end, we designed a survey for real-world developers, focusing on the key obstacles and design choices faced by practitioners in UAV-LLM integration. The analysis of 52 real-world responses shows that the integration of LLMs with UAVs remains at an early stage. Only 19.05% of teams have deployed such projects, while 59.6% have not attempted integration, citing reasons such as existing technology has met the project requirements (54.8%), insufficient performance of current LLMs (51.6%), and security risks of large language models (45.2%). The assessed technical maturity of all tasks is below three on a 5-point scale, with relatively higher scores for Information Summarization and Report Generation and LLM-based Chatbot, but lower scores for Multi-UAV Task Reasoning and Planning and Control Logic and Command Generation. Furthermore, 82.7% of practitioners prefer Onboard Control (Internal) Modes over direct control, reflecting caution toward applying LLMs in safety-critical operations. These findings suggest that while academic research emphasizes expanding task capabilities, industry adoption is constrained by performance, reliability, and safety requirements, underscoring the need for closer alignment between research and practical application.
In the following, we summarize the main contributions of this study as follows:
  • We identified and summarized nine types of tasks undertaken by LLMs in UAV systems and compared their distribution in academic literature and industrial projects. The results reveal the current applications and differences of UAV-LLM integration in academia and industry, providing a reference for future application-oriented research.
  • Through a practitioner survey, we analyzed their views on UAV-LLM integration. The findings reveal the potential reasons for the gap between academic and industrial efforts and emphasize the need to narrow this gap.
  • We summarized challenges for future development, provided recommendations for researchers and practitioners, and released our experimental data and results at https://github.com/UAV LLMs/OnlineSupplementMaterial to support future research.
[…]
2. Background
2.1. Unmanned Aerial Vehicles
As an aerial mobile platform with perception, decision-making, and execution capabilities, UAVs have been widely used in mission scenarios such as agricultural inspection (Tokekar et al., 2016), disaster response (Erdemir et al., 2020), logistics distribution (Stolaroff et al., 2018), and urban monitoring (Goodchild and Toy, 2018). With the diversification of application tasks, the dynamic operation environment, and the complexity of mission requirements, UAV systems are facing a growing demand for intelligence, including more sophisticated control strategies, more efficient path planning, and more natural human-computer interaction mechanisms (Shakhatreh et al., 2019).
To support this evolution, open-source autopilot platforms and communication frameworks have become an important foundation of UAV research and development. PX4 (Meier et al., 2015) provides a widely used flight control software stack that supports both research and industrial UAV applications. MAVLink (Koubaa et al., 2019) offers a lightweight communication protocol that allows reliable message exchange between UAVs and ground stations. The ArduPilot project (Baldi et al., 2022), with extensive documentation and community support through the ArduPilot Wiki, enables developers to build and customize UAV systems for diverse mission profiles. Traditional rule-based control methods have gradually exposed their limitations in dealing with these challenges, driving the need for UAV technology to evolve to a higher level of intelligence.
2.2. Large Language Models
Large language models (LLMs) have demonstrated strong generalization and semantic understanding capabilities in natural language processing, code generation, automatic question answering, reasoning planning, and other tasks (Naveed et al., 2023). Current mainstream LLMs (such as the GPT series, Claude, etc.) are based on the Transformer architecture. They can extract deep semantics from natural language, perform logical reasoning, and model tasks. Furthermore, LLMs have gradually acquired multimodal input processing capabilities, showing broad prospects for integration with embedded systems in the fields of visual language understanding, robot control, and task planning.
2.3. Applications of LLMs in UAV Systems
The integration of UAVs with LLMs enables flight platforms to directly understand human instructions and carry them out, lowering the barrier to use and enabling more flexible operation. Some recent studies have already explored this possibility. For example, UAV-GPT (Koubaa et al., 2025) shows how a natural command like “please patrol this area” can be translated into structured parameters that the UAV then executes. In another direction, Grounded-SAM combined with GPT-4V (Ren et al., 2024; Limberg et al., 2024) demonstrates how language models can work with visual input to label aerial images, recognize targets, and generate natural descriptions.
These practices highlight the potential of combining LLMs with UAVs. However, the current progress remains fragmented, and there is no unified framework to guide the integration of LLMs into UAV systems. This motivates our study, which aims to provide a systematic investigation into the functional roles, technical challenges, and integration patterns of UAV-LLM systems.
2.4. Card Sorting Method
In order to functionally classify and structure the literature on UAVs and large language models, this paper adopts the card sorting method, which is a cognitive organization method widely used in information architecture design and classification system construction  (Wood and Wood, 2008; Spencer, 2009). In the card sorting process, researchers first create a card for each analysis object (such as a document), which contains key information related to the classification task, such as functional description, model type, or task instance. Then, researchers classify the cards with semantic associations in a bottom-up manner, thereby gradually building a logically consistent classification system. According to the degree of presetting of categories, the card sorting method can be divided into three types: open card sorting, closed card sorting, and hybrid card sorting. In closed card sorting, all categories are preset before sorting, and users only need to classify items into fixed categories. In open card sorting, categories are not preset, but are dynamically generated by participants according to data characteristics during the sorting process. Hybrid card sorting combines the characteristics of both, allowing expansion based on existing categories to achieve both systematicity and flexibility.
[…]
3. RQ1: What is the current research status and focus of large language models in the field of UAVs?
LLMs have powerful instruction understanding, task reasoning, and interactive control capabilities. Introducing LLMs into UAV systems is expected to achieve stronger intelligent perception and task coordination capabilities. However, most existing studies focus on specific models or application cases, and there is still a lack of systematic classification of the UAV-LLM integration. This not only limits the generalization application of new technologies, but also hinders the exploration of cross-domain solutions to key problems (such as task planning, information interaction, etc.). Therefore, it is necessary to comprehensively identify and summarize the current application fields and task categories through a systematic literature review, and identify future research directions in the UAV-LLM integration.
3.1. Systematic Literature Review (SLR)
This study followed the systematic literature review (SLR) method guidelines proposed by Kitchenham et al. (Kitchenham et al., 2009) and carried out the research work according to its process. The specific process is shown in Figure 1, which includes three stages: literature search, literature selection, and data analysis. First, a literature search was conducted in databases including Google Scholar, ACM, and IEEE using a keyword-based strategy, and a total of 997 initial documents were obtained. Secondly, based on five exclusion criteria and combined with manual abstract analysis and snowball strategy (Bodin et al., 2019), 74 documents that were relevant to the research question were screened out. Finally, the selected documents were systematically read, and through card classification and inductive analysis, the answer to research question RQ1 was finally formed.
Figure 1. Methodological process for answering RQ1 via systematic literature review.
3.1.1. Literature Search
To collect high-quality research literature on the theme of UAV-LLM integration, we first conducted a preliminary search using the Google Scholar engine and obtained 997 candidate papers. In our data collection strategy, we required that the paper title contain at least one of the keywords “Unmanned Aerial Vehicle (UAV)” or “Large Language Model (LLM)” to ensure broad coverage of the field. However, many works primarily discuss other aspects of UAV systems, such as hardware platform design, flight control engineering, or photogrammetry applications (Gupta et al., 2015; Shakhatreh et al., 2019; Colomina and Molina, 2014), which fall outside the scope of our focus on UAV-LLM integration. Therefore, we further filtered the corpus by requiring that the abstracts explicitly mention both “UAV” and “LLM”, ensuring that the final dataset directly relates to the integration of these two technologies.
3.1.2. Literature Selection
Although all 997 papers retrieved in our search contained the keywords “Unmanned Aerial Vehicle (UAV)” and “Large Language Model (LLM)” in their abstracts, some of these papers did not actually discuss the integration of both technologies. For example, some studies focused on UAV systems but emphasized hardware platforms or flight control design, and were mistakenly included in the initial screening due to mentioning the technical background of LLM (Mueller, 2007). To improve the accuracy of the screening, we eliminated irrelevant articles based on the following five exclusion criteria:
  1. (1)
    Papers not written in English;
  2. (2)
    Master’s or doctoral dissertations;
  3. (3)
    Research topics that are generalized and do not focus on the combination of UAV and LLM technology;
  4. (4)
    Research that is not related to large language models;
  5. (5)
    Research that is not related to UAV technology.
To reduce bias in the filtering process, we employed a closed-card classification method to assess the relevance of the papers systematically. The classification criteria were “relevant” and “irrelevant”. The filtering was carried out independently by two graduate students, who expanded the initial dataset using snowball sampling (Bodin et al., 2019) and reviewed the abstracts of all papers. In cases of inconsistent initial judgments, the two researchers discussed and reached a consensus. After multiple rounds of filtering, we selected 74 relevant and representative papers from the initial 997 for further task categorization and quantitative analysis.
3.1.3. Data Analysis
After completing the literature filtering, we conducted an in-depth analysis of the 74 selected papers and systematically categorized the tasks described in them. Given the objective of this study to explore the practical tasks of LLMs in UAV systems and the lack of a unified classification system, we chose an open-card classification method to identify potential, previously unknown task categories. Specifically, two researchers independently read each paper and extracted specific applications of LLMs in UAV systems. During the data extraction, we first identified the raw task expressions, such as “natural language parsing of flight commands” and “dialog-based multi-round path replanning”, and then condensed these into task intents with standardized naming, such as “task parsing” and “path reasoning and planning”. Subsequently, we grouped similar tasks into higher-level categories based on their semantic relevance and technical implementation logic, ultimately resulting in nine task types, which were further abstracted into four core stages of the UAV workflow. The entire card classification and clustering process was iteratively refined through multiple discussions, ensuring consistency and reliability in the classification system.
3.2. Results
Based on analysis results, we have summarized the UAV workflow into the following four steps: Information Input, Task Planning, Control Execution, and Result Feedback, which are shown in Figure 2. Each step is associated with specific tasks that LLMs assist with. In the Information Input phase, LLMs support Multimodal Information Parsing and Natural Language Command Parsing. In the Task Planning phase, LLMs are involved in Single-UAV and Multi-UAV Task Reasoning and Planning, as well as RAG-based Decision Support. The Control Execution phase utilizes Control Logic and Command Generation/Optimization. Finally, the Result Feedback phase incorporates Information Summarization and Report Generation and the development of LLM-based Chatbot.
Figure 2. Illustration of the UAV workflow and LLM task modules.
The analysis of 74 research papers enabled us to identify nine tasks that LLMs currently undertake within UAV systems; the detailed characteristics of each task are summarized in Table 1. These nine tasks cover the entire workflow from information input to task planning, control execution, and result feedback. Specifically, they include: Natural Language Command Parsing (25/74,33.78%), Multimodal Information Parsing (23/74,31.08%), Single-UAV Task Reasoning and Planning (29/74,39.19%), Multi-UAV Task Reasoning and Planning (11/74,14.86%), RAG-based Decision Support (15/74,20.27%), Control Logic and Command Generation (27/74,36.48%), Control Logic and Command Optimization (9/74,12.16%), Information Summarization and Report Generation (19/74,25.68%), and LLM-based Chatbot (15/74,20.27%). These task modules, as part of the intelligent UAV system workflow, can be flexibly selected based on specific scene requirements, supporting autonomous perception, intelligent decision-making, and multi-UAV collaboration capabilities in complex task environments. It should be noted that a single paper may involve multiple tasks, so the summed proportions exceed 100%.
Table 1. Tasks Performed by LLMs in UAV Systems.
UAV Workflow Task Name Description Example Application
Information Input Natural Language Command Parsing Converts user natural-language commands into structured task intents. “Please fly over the building complex and take photos” $\rightarrow$ {fly, to: coords, action: photo}
Multimodal Information Parsing Analyses and annotates UAV imagery, video, and remote-sensing data. Image $\rightarrow$ textual description, bounding-box targets, environmental semantic labels.
Task Planning Single-UAV Task Reasoning and Planning Performs multi-stage task decomposition and environmental path planning through reasoning. Detect weather-affected routes and reschedule task order accordingly.
Multi-UAV Task Reasoning and Planning Allocates tasks among multiple UAVs, dynamically adjusts plans, and executes cooperative control. Assign tasks to five UAVs and dynamically switch targets.
RAG-based Decision Support Retrieves regulations, maps, and domain knowledge to enhance decision-making. Determine whether an area is a no-fly zone / find the energy-optimal route.
Control Execution Logic and Command Generation Outputs low-level control code (e.g., PX4, MAVLink) according to the mission plan. Automatically generate flight-control scripts and inject them into the controller.
Control Logic and Command Optimization Senses the spectrum environment and network topology to optimise communication strategies and ad-hoc networking. Select interference-free channels; rebuild relay networks in disaster zones.
Result Feedback Information Summarization and Report Generation Extracts key points from UAV status, logs, and sensor data to produce concise summaries or reports. Generate a daily disaster-monitoring report from the logs of multiple UAVs.
LLM-based Chatbot Manages tasks, status queries, and user feedback through question-and-answer dialogue. “Have you completed the target?” $\rightarrow$ “The task in Area 2 is complete.”
Finding 1: Academic research is distributed across the four main areas of UAV workflows: Information Input, Task Planning, Control Execution, and Result Feedback. Within these workflows, the four most common tasks are: Single-UAV Task Reasoning and Planning (39.19%), Logic and Command Generation (36.48%), Multimodal Information Parsing (31.08%), Natural Language Command Parsing (33.78%).
3.2.1. Information Input
The Information Input module is the starting point of the UAV workflow, primarily responsible for converting external commands and perception data into task information that the system can understand and execute (Colomina and Molina, 2014; Zhang et al., 2023). The tasks performed by large models in this module include the parsing of natural language commands and the analysis of multimodal data (such as images, videos, and remote sensing data), ensuring that the system can accurately understand and respond to task requirements.
Natural Language Command Parsing Natural Language Command Parsing refers to the use of large models’ natural language understanding capabilities to convert user-inputted natural language commands into structured task information. In UAV systems, before the use of LLMs, this task mainly relied on traditional speech recognition and rule-based engine technologies (Gupta et al., 2016), which often lacked sufficient understanding capabilities and could not fully analyze the context and implicit intentions of commands, leading to lower flexibility and accuracy in task execution. With LLMs, the system can accurately parse complex instructions and convert them into executable task commands through their powerful natural language understanding ability (Ahn et al., 2022). This enables UAV systems to process complex commands more intelligently, improving task execution precision and efficiency. For example, when a user inputs “Fly over the west side of the bridge and take a photo”, the system can accurately parse the flight action, target location, and task type, and generate specific execution commands. This technology has been successfully applied in systems such as Swarm-GPT (Jiao et al., 2023), TypeFly (Chen et al., 2023), and AutoHMA-LLM (Yang et al., 2025), greatly enhancing the intelligence of UAV task execution.
Multimodal Information Parsing Multimodal Information Parsing refers to the ability of large models to process non-linguistic inputs and understand scene semantics, enabling the parsing of UAV multimodal data to provide essential contextual understanding and support for task planning. In UAV systems, traditional methods relied on single sensory channels, often limiting their effectiveness in handling complex environments or multiple data sources, which led to an incomplete understanding of environmental information (Zhu et al., 2021). With the introduction of LLMs, the system can combine visual encoders (e.g., BLIP-2 (Li et al., 2023), SAM (Kirillov et al., 2023), Grounding DINO (Liu et al., 2023)) and form complete vision-language (V-L) models (Zeng et al., 2023b), allowing it to parse images, videos, and remote sensing data from UAV sensors like cameras and radar, extracting semantic information and generating language descriptions. This path significantly enhances the UAV’s ability to perceive and understand complex environments, resulting in improved performance in remote sensing, inspection, and security monitoring tasks. Representative systems such as AeroReformer (Li and Zhao, 2025), CLIPSwarm (Pueyo et al., 2024), and video anomaly detection (Zhao et al., 2024b) demonstrate the powerful capabilities of large models in multimodal data parsing, greatly enhancing UAV perception and decision-making support.
Finding 2: The highest attention in the nine tasks executed by LLMs in UAV systems is 39.19%, while the lowest is 12.16%. This difference suggests that academic attention is more evenly distributed across various tasks, with no single area of high attention.
3.2.2. Task Planning
Task Planning is primarily responsible for transforming task objectives into executable operation paths, ensuring the efficient and accurate execution of tasks. By optimizing task reasoning, path planning, and task allocation, the task planning module can handle various task requirements and complex environmental changes (Shakhatreh et al., 2019). This module involves not only the execution of single-UAV tasks but also the management of multi-UAV collaborative tasks and decision support based on external knowledge.
Single-UAV Task Reasoning and Planning Single-UAV Task Reasoning and Planning involves using large models to reason through multi-stage tasks performed by a single UAV, generating optimal paths and task execution plans. Before LLMs, this task typically relied on traditional rule-based engines or simple path-planning algorithms (Zhu et al., 2019), which struggled to adapt to dynamic changes in the environment, task priorities, and external constraints, resulting in lower task execution efficiency and flexibility. With LLMs, the system can perform real-time reasoning based on task planning, dynamically adjusting task execution plans and path planning to ensure tasks are completed in complex environments (Huang et al., 2024). This technology path enables the system to automatically adjust task order and paths based on real-time changes such as weather and obstacles. For example, when a path is affected by weather, the system will rearrange the task sequence to avoid risks. NavAgent (Liu et al., 2024) combines graph neural networks with GPT-4 for outdoor landmark-based navigation path reasoning. Additionally, reinforcement learning integration methods (Xiang et al., 2024) and the TAMP system (Sautenkov et al., 2025) further demonstrate this capability, enhancing task execution intelligence and adaptability.
Multi-UAV Task Reasoning and Planning Multi-UAV Task Reasoning and Planning refers to using large models to solve the problem of how multiple UAVs can collaborate and complete tasks. Traditional multi-UAV task scheduling relied on predefined rules and simple task allocation methods (Chung et al., 2018), which often struggled to handle dynamic environments and real-time changes, resulting in low collaboration efficiency. With LLMs, the system can intelligently allocate tasks and dynamically adjust task assignments and path planning across multiple UAVs, improving collaboration efficiency and flexibility. This technology path allows the system to adjust task areas and UAV allocations in real-time based on task priorities, environmental changes, and target status. For example, during monitoring tasks, the system automatically switches task areas based on target status, enabling efficient multi-UAV collaboration. AutoHMA-LLM (Yang et al., 2025) achieves multi-level, multi-agent collaboration, with LLM supporting dynamic task decomposition and real-time adjustments. Swarm-GPT (Jiao et al., 2023) and FlockGPT (Lykov et al., 2024) also demonstrate the effective application of this technology, significantly improving the efficiency and reliability of multi-UAV cooperative tasks.
RAG-based Decision Support RAG-based Decision Support refers to using large models combined with external knowledge to enhance the decision-making ability of UAVs. Before the use of LLMs, traditional decision support systems relied on static rule bases and limited external data, which could not flexibly handle uncertainty in complex environments. With LLMs, the system uses Retrieval-Augmented Generation (RAG) technology (Lewis et al., 2020) to quickly retrieve and analyze external data such as maps, no-fly zone information, and regulatory databases, providing precise decision support for task planning, ensuring compliance, safety, and efficiency. This technology path allows the system to assess risks in real-time during task execution and make optimal decisions. For example, SAGIN network management (Tang et al., 2024) and regulatory determination papers (Liu et al., 2024; Javaid et al., 2024) demonstrate knowledge-enhanced intelligent decision-making, improving the system’s ability to make decisions in complex environments and ensuring the successful completion of tasks.
3.2.3. Control Execution
The Control Execution module is responsible for converting the outputs of the task planning phase into specific flight control commands, ensuring that the UAV executes tasks in accordance with the requirements (Shakhatreh et al., 2019). At this stage, the system not only generates control logic and commands but also involves real-time optimization and adjustment of control strategies to cope with complex and dynamic environmental conditions.
Control Logic and Command Generation Control Logic and Command Generation refers to the use of LLMs to transform task planning results into executable flight control commands for UAVs (e.g., PX4, MAVLink), guiding UAVs to perform flight tasks. Before the introduction of LLMs, this task typically relied on manually written control logic or predefined command sets, resulting in a strong dependence on specialized programming knowledge and low efficiency in control logic generation (Koubaa et al., 2019; MAVLink Core Team, 2024). With the introduction of LLMs, the system can automatically generate flight control logic, ensuring that tasks are executed efficiently along the predetermined path. This technology path enhances the automation of control command generation, reducing human intervention and development time. For example, TypeFly (Chen et al., 2023), PX4 command generation-related papers (TAZIR et al., 2023; Zhu et al., 2024) demonstrate the automatic generation of control logic and code, effectively reducing the reliance on professional programming knowledge.
Control Logic and Command Optimization Control Logic and Command Optimization involves real-time analysis of environmental data and task feedback to optimize control strategies. Traditional methods typically rely on fixed control strategies or simple adjustment rules, which are limited in complex or dynamically changing environments (Campion et al., 2018). With LLMs, the system can assist in optimizing UAV communication strategies, link configurations, and anti-jamming designs, improving communication efficiency and stability within the swarm (Wei et al., 2025; Zhou et al., 2024). This technology path enables real-time adjustment of control logic in complex or uncertain environments, ensuring that UAVs can perform tasks efficiently and stably, especially when facing flight interference, communication issues, or other unforeseen circumstances. For example, SAGIN network management (Tang et al., 2024) demonstrates the application of LLMs in cross-layer communication network scheduling and security protection, solving resource allocation problems in complex environments. Papers (Li et al., 2025b; Yaqoot et al., 2025; Zhu et al., 2025) further provide specific implementations of this task, enhancing the effectiveness of control logic and command optimization.
3.2.4. Result Feedback
The Result Feedback module is responsible for providing information output and feedback after task execution, including task status summaries, report generation, and interaction feedback with users. This module supports real-time communication between the UAV system and the user, ensuring transparency in task execution and efficient feedback (Wang et al., 2024b; Zhang et al., 2024).
Information Summarization and Report Generation Information Summarization and Report Generation refers to the task where large models analyze UAV status, logs, perception data, and flight trajectories to generate concise and clear task reports automatically. Before the introduction of LLMs, reports were typically manually generated or relied on traditional analytical methods, which were often inefficient and lacked the flexibility to adapt to dynamic task requirements quickly. After integrating LLMs, the system can automatically generate summaries of task progress, key events, and potential risks, helping users quickly understand the task status and make timely decisions. This approach significantly improves the automation and real-time nature of task reporting. For example, after completing a monitoring task, the system can generate daily reports that include task execution status and environmental changes, enhancing the readability of task information and decision-making efficiency. Public attitude analysis (Zhu, 2024) and the Hawkeye video anomaly detection system (Zhao et al., 2024b) use LLMs to provide real-time summaries of complex data, assisting decision-makers in quickly understanding task status and risks.
LLM-based Chatbot An LLM-based Chatbot refers to the use of large models to enable natural language dialogue between humans and the UAV system. Traditional task queries and feedback were often handled through fixed interfaces or simple command responses, which lacked personalized and flexible user experiences. With the introduction of LLMs, the system can achieve task inquiries, progress feedback, and status updates through multi-turn dialogues, enhancing the intelligence and personalization of the interaction between the UAV system and the user (Zhang et al., 2024). Through this approach, users can interact more naturally with the system and quickly obtain task status updates to make real-time decisions. For example, users can ask, “How much of the current task is completed?" or "Are there any abnormalities?” The system can quickly and accurately respond, helping users stay updated on task progress. Representative UAV systems have integrated conversational interfaces for querying execution status, generating explanations, and issuing prompts (Zhao et al., 2024a; Aikins et al., 2024); in addition, mature voice and multimodal speech front-ends for UAV control can be combined with LLM dialogue further to enhance usability (Park et al., 2023; Choutri et al., 2022; Oneaţă and Cucu, 2021).
[…]
4. RQ2: What are the differences between academia and industry in integrating UAVs with large language models?
To assess the current level of industrial development in integrating LLMs with UAVs and to identify the main challenges, we compiled a dataset of real-world projects that combine these technologies. We then compared and analysed the collected papers and projects, using the four core application domains and nine task categories defined earlier, in order to clarify the gaps between academic research and industrial practice and to explore the reasons for those gaps.
4.1. Data Collection and Processing
In the data collection process, we surveyed GitHub to identify industry projects integrating LLMs with UAVs. Specifically, we conducted keyword-based retrieval using two curated sets of keywords covering common terms related to UAVs (e.g., “drone”, “UAV”, “quadcopter”) and LLMs (e.g., “GPT”, “large language model”, “LLM”). We retained repositories where at least one UAV-related keyword and one LLM-related keyword co-occurred in the title, README, or description. Subsequently, two authors independently reviewed each candidate repository to verify its relevance. Ultimately, we identified 56 projects that provided either complete engineering solutions or prototype demonstrations of UAV-LLM integration.
To investigate the gap between academia and industry in the integration of LLMs with UAVs, we conducted a comparative analysis based on the task classification framework of LLM applications in UAVs established in RQ1. The academic data source consists of 74 relevant publications obtained through a systematic literature review (SLR) conducted in RQ1, while the industry data source comprises 56 real-world projects collected by two researchers from major technical and social platforms. The task classification framework was developed by the two researchers in RQ1 using an open card sorting method, which distilled nine types of tasks performed by LLMs in UAV scenarios and five core application domains. We annotated and categorized each data point in both datasets according to the tasks and application directions of LLMs, mapping them to the four core domains and nine specific tasks shown in Figure 1. The frequency distribution of each category within the academic and industrial ecosystems was then analyzed. This entire process involved multiple rounds of iteration and discussion between the two researchers to reach consensus, ensuring scientific rigor and consistency in the results. The outcome provides a solid foundation for identifying potential commonalities and differences between academia and industry in this emerging field.
Figure 3. Statistics on the application of LLMs in UAVs in academia and industry.
4.2. Results
4.2.1. Current State of Industry
Firstly, we conducted a statistical analysis of the tasks performed by LLMs in UAV systems for each industrial project. The results include: Natural Language Command Parsing (29/56, 51.79%), Multimodal Information Parsing (17/56, 30.36%), single-UAV Task Reasoning and Planning (28/56, 50%), Multi-UAV Task Reasoning and Planning (8/56, 14.29%), RAG-based Decision Support (5/56, 8.92%), Control Logic and Command Generation (40/56, 71.43%), Control Logic and Command Optimization (2/56, 3.57%), Information Summarization and Report Generation (12/56, 21.43%), and LLM-based Chatbot (8/56, 14.29%). We observed that developers are most focused on flight control in UAV systems, with projects performing the task of Control Logic and Command Generation accounting for 71.43% of all projects. Natural Language Command Parsing and Single-UAV Task Reasoning and Planning are also popular, accounting for 51.79% and 50%, respectively, indicating an increasing demand in the industry for task automation and intelligent interaction.
Finding 3: In over 50% of the projects, the tasks handled by large models are concentrated in areas such as flight control, task planning, and human-machine interaction, with less than 10% of the projects focusing on other tasks.
In contrast, tasks such as RAG-based Decision Support and Control Logic and Command Optimization are less commonly applied, which could be because these tasks require more external knowledge bases and datasets. Additionally, Multimodal Information Parsing and Multi-UAV Task Reasoning and Planning have seen some applications in the industry but still face challenges related to real-time performance, computational resource demands, and the inherent complexity of these tasks.
4.2.2. The differences between academia and industry
We analyzed the literature and projects that integrate UAVs with LLMs, presenting their specifics in Figure 3 and identifying three key differences between academia and industry in the integration of UAVs with LLMs. In Figure. 3, “UAV Workflow” and “Tasks performed by LLMs in UAVs” represent the four UAV workflows and the nine tasks performed by LLMs in UAV systems that we identified in RQ1. “(Academia) Literature ID” and “(Industry) Project ID” refer to the IDs of the literature and projects we collected. We have analyzed the UAV tasks involved in each of the literature and projects.
Prevalence and Distribution of LLM-UAV projects. In terms of prevalence, there are currently fewer industrial projects combining large language models with UAVs than academic research, which could be attributed to the early development stage and high development cost of LLM-UAV integration. The focus range of academic research on the nine tasks spans from 12.16% to 39.19%, while in industry, the range is from 3.57% to 71.43%. In terms of scope and depth, this difference indicates that academic research has a more dispersed focus across various tasks, with a broad research scope and no apparent concentration of high attention on specific tasks. Instead, it covers multiple key areas such as multi-UAV collaboration, decision support, and environmental perception. In contrast, industrial projects exhibit a trend of task centralization, with higher attention given to tasks directly related to product interaction and execution, focusing more on practicality and the automation of task execution.
Application Fields. Academic research focuses on cutting-edge areas, emphasizing theoretical modeling and task optimization, including algorithm optimization, data fusion, and the application of external knowledge. For example, academic research addresses multi-UAV collaboration and decision support using RAG and reinforcement learning techniques (Liu et al., 2024; Javaid et al., 2024), combines Q-Learning, graph neural networks, and LLM reasoning to achieve dual optimization of communication and task execution (Li et al., 2025b), and utilizes LLMs for link quality prediction and dynamic resource allocation (Tang et al., 2024). On the other hand, the industrial sector faces higher real-time requirements and emphasizes practical operability and scalability, focusing on how to improve task execution efficiency and control accuracy in real-world applications. For example, in more than 50% of the projects, large models are responsible for tasks such as Natural Language Command Parsing, Control Logic and Command Generation, and Single-UAV Task Reasoning and Planning. These tasks are concentrated in areas like flight control, task planning, and human-machine interaction.
Finding 4: Regarding the development of RAG-based Decision Support and Control Logic and Command Optimization, academic research is primarily focused on these areas, with respective proportions of 20.27% and 12.16%. In contrast, their application in industrial projects is relatively limited, with respective proportions of 8.92% and 3.57%.
Different Integration Modes Between LLMs and UAVs. To gain a deeper understanding of the technical details of LLMs’ application in the UAV field in both academia and industry, we identified three main integration modes of LLMs with UAVs based on the location of the LLM and the interaction between the LLM and UAV. These modes are: External (Offboard Invocation) mode, Hybrid (Onboard Assistance) mode, and Internal (Onboard Control) mode. We researched their distribution in current academic studies and industrial projects.
  1. (1)
    Offboard Invocation (External): In this mode, the LLM is deployed outside the UAV, either in the cloud or at a ground-based edge computing node. It receives task information from the UAV via remote APIs and returns task suggestions and semantic interpretations. This mode offers flexible deployment and strong computational power but suffers from high communication latency and limited real-time performance, making it more suitable for non-real-time tasks such as large-scale inference and complex semantic parsing. The use of this mode in academia (52.7%) is higher than in industry (41.07%), reflecting that academic research is still largely focused on cloud-based invocation and experimental validation, with tasks concentrated on algorithm verification and offline task planning. In contrast, industry has reduced the proportion of offboard invocation, indicating that some tasks have migrated to edge/onboard systems to reduce latency and improve real-time performance.
  2. (2)
    Onboard Assistance (Hybrid): In this mode, the LLM is deployed on an edge computing device within the UAV (such as Jetson, RK3588, etc.). It does not directly generate control commands but acts as a decision-support module, collaborating with the UAV’s flight control system to support task planning, path inference, and multi-UAV coordination. This mode balances intelligence and stability, making it suitable for complex tasks such as real-time perception, path adjustment, and local decision-making. The usage of this mode in industry (50%) is higher than in academia (38%), suggesting that in real-world industrial deployments, real-time performance and stability are considered more important than extreme inference capabilities. Academic research, however, may be limited by hardware constraints and deployment complexity.
  3. (3)
    Onboard Control (Internal): In this mode, the LLM is deeply integrated into the UAV’s flight control system, directly responsible for task decomposition, action decision-making, and control command generation. Its advantage is that the UAV’s flight control directly relies on the LLM, providing good real-time performance. It also offers high real-time capability and autonomy, making it a key architecture for achieving fully autonomous flight. However, the disadvantage is the high demand for local computational resources, which is limited by the model size. The usage of this mode is low in both academia (5.4%) and industry (10.71%), reflecting a cautious attitude. This mode is currently more commonly used in high-value, specific tasks with clear ROI (such as rescue, defense, and mining). It can be speculated that, as technology advances, particularly with edge computing and lightweight models, this mode will have significant growth potential in the future.
Finding 5: In the three integration methods of LLMs and UAVs, academia most commonly adopts External-offboard invocation(52.7%), while industry tends to prioritize Hybrid-onboard assistance (50%). This indicates that academia focuses more on algorithm validation and performance evaluation in controlled environments, whereas industry places greater emphasis on real-time performance and the practical application of resource constraints. .
Finding 6: Currently, LLMs are rarely directly involved in UAV control, but the development of edge computing and lightweight models may bring new possibilities for the application of large models in UAVs.
[…]
5. RQ3: How do real-world practitioners perceive these differences, and what factors contribute to them?
Section 4 presents preliminary findings that reveal the differences in UAV plus LLM integration between academia and industry. To further explore the potential reasons behind these differences and identify possible solutions to bridge the gap, we conducted an online survey targeting real-world UAV developers. The goal was to gain deeper insights into practitioners’ perspectives on integrating LLMs into UAV systems, as well as to examine the key obstacles and design choices encountered in this integration process.
5.1. Survey Design
The UAV survey is divided into four sections.
Demographics. (Q1-Q4) This section collected participants’ professional background to ensure relevance and interpretability of their responses. Participants were asked about their primary domain (e.g., UAVs, robotics, LLMs) (Q1), responsibility (hardware/software/algorithm development, testing, research, etc.) (Q2), years of experience (Q3), and geographical location (Q4).
Practices and Challenges. (Q5-Q6) We explored whether the practitioners had ongoing projects combining LLMs and UAVs (Q5), and if not, what challenges (e.g., cost, technical limitations, regulation, unclear business models) had prevented integration (Q6).
Application Maturity and Integration Approach. (Q7-Q10) Participants were asked to rate the maturity of various LLM application scenarios in UAVs, such as Natural Language Command Parsing, Multimodal Information Parsing, Single-UAV Task Reasoning and Planning, and Multi-UAV Task Reasoning and Planning, as well as other related tasks (Q7). Additional open-ended questions inquired about other potential applications desired by practitioners (Q8), the implementation approaches of UAV-LLM integration(e.g., External, Hybrid, and Internal) (Q9), and the LLM’s capabilities desired for UAV tasks (Q10).
Industry-Academia Comparison. (Q11-Q13) To understand the alignment between research and practice, we asked whether participants referred to academic work in their projects, and collected their views on the differences between academic and industrial efforts (Q11). Specific factors such as development lifecycles, technical requirements (Q12), and integration strategies were covered. Furthermore, we investigated the perceived underlying reasons for these differences, such as real-world requirements for technical maturity, performance, safety, and cost (Q13).
5.1.1. Participant Recruitment.
Our target practitioners for the survey were UAV developers. Our goal was to distribute the survey to GitHub developers who had contributed to UAV-related projects. First, we searched for UAV-related repositories on GitHub using keywords such as “UAV”. Then, we used the GitHub Search API to collect the email addresses and names of contributors to the selected repositories. In total, we obtained 802 developer email addresses and sent out invitations to participate in the survey. Additionally, we shared the survey with several industry collaborators working at well-known companies, such as China Mobile Communications Corporation, with a total of six industry partners involved. Since some developers might not be familiar with the term “UAV,” we provided a brief explanation of the concept in the email to avoid confusion.
5.1.2. Data Analysis.
We received a total of 55 valid responses from 15 different countries, with a response rate of approximately 6.85%. Among these practitioners, we excluded three individuals who indicated that they had no development experience related to UAVs, robotics, or LLMs, and conducted further analysis on the remaining 52 participants. The average experience of the remaining practitioners was 5.9 years. Their primary responsibilities included hardware development (12, 23.01%), software development (37, 71.15%), algorithm development (25, 48.08%), testing (9, 17.31%), and research (13, 25%). Note that a practitioner may concurrently hold multiple roles, such as developing and testing software.
5.2. Results
5.2.1. Practices and Challenges in UAV-LLM integration
Real-world Practices. In the survey, 40.4% of the practitioners have started to advance projects that integrate LLMs with UAVs), with 19.05% of them having fully deployed the technology, and 80.95% of them being in the proof-of-concept phase. This indicates that, although the integration of LLMs with UAVs is still in the exploratory stage, some practitioners have already successfully implemented the initial application of this technology. However, 59.6% of the practitioners have not yet initiated this project. For those practitioners who have not implemented the project, we further inquired about the specific reasons for not advancing this technology. The survey revealed the main challenges.
Challenges Perceived by Practitioners. The analysis shows that the obstacles mentioned by practitioners are mainly related to technology, resources, and market factors. In particular, regarding technology, 54.8% of the practitioners believe that existing technologies already meet their needs, and therefore, they are not in a rush to introduce new technological solutions. This suggests that in some traditional UAV application scenarios, the existing technology and control systems are mature enough to reliably complete tasks without the need for more complex LLM technology. 51.6% of practitioners believe that existing large models cannot meet performance requirements, reflecting the current shortcomings of LLMs in terms of performance, computational power, and real-time capabilities, which are still insufficient to meet the high-performance demands in the UAV field. Additionally, 45.2% of practitioners pointed out the security risks of large models, with some noting that “LLMs are not being used on board in safety-critical applications” and “LLMs should not be on a critical system due to their ‘black box’ nature.” This indicates that data privacy, as well as the transparency and interpretability of models, remain major obstacles to the widespread application of LLMs in the UAV domain.
In addition to technological issues, 25.8% of practitioners mentioned the high costs, noting that the high computational resource demands and related hardware limitations make it difficult for these practitioners to accept when considering cost-effectiveness. 22.6% of practitioners stated that the lack of training data or fine-tuning methods limits their application, preventing them from effectively utilizing the existing large models. Furthermore, 29.0% of practitioners believe that the lack of relevant knowledge and technology is another reason for not advancing this project.
Finding 7: Currently, 40.4% of practitioners have attempted to integrate LLMs with UAVs in their projects. The main reasons for not attempting are “Existing technology has met the project requirements” (54.8%), “Existing large language models cannot meet performance requirements” (51.6%), and “Security risks of large language models” (45.2%).
Finding 8: Some practitioners also expressed doubts about the application of LLMs in critical systems such as UAVs. This suggests that the integration of LLMs with UAVs is still in the early stages of exploration. Although some practitioners have already started applying them, there is no urgent demand from the industry, and they face significant technical obstacles. To advance this field, it is necessary to enhance the performance, scalability, and reliability of LLMs in critical tasks.
Table 2. Practitioners’ perceptions on tasks maturity
[…]
Natural Language Command Parsing 2.89 57.7%
Multimodal Information Parsing 2.88 51.9%
Single-UAV Task Reasoning and Planning 2.77 48.1%
Multi-UAV Task Reasoning and Planning 2.40 38.5%
RAG-based Decision Support 2.88 55.8%
Control Logic and Command Generation 2.51 40.4%
Control Logic and Command Optimization 2.59 44.2%
Information Summary and Report Generation 2.96 59.6%
LLM-based Chatbot 2.96 61.5%
Application Maturity Perceived by Practitioners. To assess the current industrial acceptance and technical maturity of large models in UAV tasks, we invited practitioners to rate the technical maturity of various tasks of LLMs in UAVs. We used the Likert scale with a scale from 1 to 5, with 1 indicating “completely immature,” 3 representing neutral, and 5 indicating “very mature”. The results, as shown in  2, indicate that the average score for the technical maturity of all tasks is below 3, indicating that while LLMs hold significant application potential in these areas, they also face numerous challenges. Among these, practitioners considered LLMs to have relatively higher technical maturity in Natural Language Command Parsing and Information Summarization and Report Generation, with scores of 2.89 and 2.96, respectively. These tasks involve basic processing of language and data, and existing LLMs can provide effective support in these areas, resulting in a relatively higher technical maturity. However, for more complex tasks, such as Multi-UAV Task Reasoning and Planning (score 2.40) and Control Logic and Command Generation (score 2.51), the technical maturity was lower. This suggests that, while multi-UAV coordination and autonomous control theoretically have significant potential, there are still many technical obstacles in effectively applying LLMs to these complex tasks in practical industrial applications. These challenges need further improvement in areas such as accuracy, real-time performance, and collaborative computing capabilities.
Additionally, two practitioners specifically highlighted potential application scenarios for LLMs in the UAV field, including: supporting users and new developers in setting up UAV systems (e.g., helping configure ArduPilot Wiki) and applications in areas such as natural disaster response, maritime accidents, and search and rescue (SAR), environmental monitoring and hazard detection, and urban safety and public monitoring. This highlights the potential for LLMs in the UAV field in practical industrial applications, particularly in high-complexity and multi-task processing scenarios, which remain to be further explored.
Finding 9: In terms of technology maturity perceived by practitioners, Natural Language Command Parsing (score 2.89) and Information Summarization and Report Generation (2.96 points) received relatively higher scores, while Multi-UAV Task Reasoning and Planning (score 2.40) and Control Logic and Command Generation (score 2.51) scored lower. This indicates that current LLMs can provide effective support for tasks involving basic language and data processing, but tasks that require higher accuracy, real-time performance, and computational resources still need further improvement to meet the decision support and real-time response requirements of existing UAVs.
Common Integration Approaches. To gain a deeper understanding of the technical details of integrating LLMs with UAVs, we explored the main integration methods of LLMs in the UAV field. Specifically, we asked practitioners to select from various integration methods. The survey results show that Hybrid-Onboard Assistance is the most common method, with 82.7% of practitioners choosing this option. In this approach, LLMs typically do not directly control the UAV’s flight but serve as a supporting tool for task planning, path reasoning, and decision support. This suggests that, currently, LLMs in the UAV field are more focused on enhancing decision-making capabilities to improve the intelligence of task execution, rather than directly intervening in the flight control system.
In comparison, External-Offboard Invocation (i.e., deploying large models on cloud or ground devices to provide decision support via external interfaces) and Internal-Onboard Control (i.e., embedding large models into the UAV’s control system to generate flight control commands directly) were selected by 67.3% and 26.9% of practitioners, respectively. The application of these two methods is relatively limited and, in practical use, is not as common as the Hybrid-Onboard Assistance approach. This reflects that the application of LLMs in the UAV field is still in the early stages, with the industry more reliant on enhancing decision support rather than direct control.
Although LLMs show strong potential in auxiliary tasks, their stability, interpretability, and real-time performance have not yet met the high standards required for industrial applications in core tasks like flight control. Therefore, LLMs are currently more applied in low-risk scenarios such as Task Planning and decision support, while their use in high-risk areas like flight control remains highly restricted. One practitioner in the survey pointed out, “I do not think LLMs are needed for tasks inboard the uav. AI can have a place on board for image processing to add sensors. But AI should never be put in direct control of a uav. It is an unnecessary, lazy, and non-robust approach to vehicle control.” This clearly reflects the cautious attitude of the industrial sector towards the use of LLMs in critical UAV control tasks, with the belief that LLMs are more suitable for supporting tasks rather than directly participating in flight control.
The Need of UAV-LLM integration. Additionally, we asked practitioners about the level of need for LLMs in different UAV tasks. The results show that practitioners have a higher demand for Information Input (score 2.42) and Task Planning (score 2.53). This indicates that LLMs have significant application potential in processing user command translation, perception data, and Task Planning. These tasks require high data processing, reasoning, and planning capabilities, and therefore, there is a strong demand for LLMs’ involvement in these areas. In contrast, the demand for Control Execution (score 2.05) is lower, suggesting that in the areas of flight control and command generation, existing technologies are already well-suited to meet the needs, or the current large language model technology cannot be fully relied upon. This indicates that in real-time flight control systems, traditional control algorithms and hardware systems still dominate.
Finding 10: Regarding task demands, Information Input (score 2.42) and Task Planning (score 2.53) received higher scores, and 82.7% of practitioners believe that Onboard Assistance is the primary way to integrate large language models with UAVs. This indicates that, at present, LLMs in the UAV domain are more focused on supportive tasks rather than directly intervening in the flight control system.
5.2.2. Industry-Academia Comparison.
Practitioner’s Perception of the Academia-Industry Gap. To understand the practical application of academic research in industrial settings, we asked practitioners whether they refer to academic research in their actual work and included an open-ended question asking for the reasons. The results show that 69.2% of practitioners indicated that they do refer to academic research. The main reasons for this are as follows: gaining cutting-edge knowledge, verifying the accuracy of information, accelerating development speed, and learning advanced algorithms and technologies. In particular, in the area of algorithm development, academic research provides developers with new ideas and methods, which have actively contributed to improving work efficiency and development quality. For example, one practitioner mentioned, “I am implementing algorithms for paths’ planning, etc., from the academic publications, I also use them during brainstorming of new ideas and to learn state-of-the-art.
Finding 11: 69.2% of practitioners stated that they would reference academic research results in their actual work to stay updated with the latest technologies, which suggests that academic research plays an important role in driving UAV-LLM practices.
Although most practitioners tend to refer to academic research, a small portion of practitioners do not directly refer to academic outcomes. In this regard, we also asked for their views on the differences between academia and industry in the process of integrating UAVs with large models. According to the survey results, 51.9% of practitioners believe that industry is progressing faster than academia. This suggests that while academia is advancing in theoretical research and exploring new technologies, the actual application and implementation in industry are happening at a faster pace. In contrast, 36.5% of practitioners believe that academia is progressing faster than industry. This reflects that, in basic theoretical research and technological innovation, academia’s breakthroughs provide new ideas and directions for the industry. However, because academic research is not always aligned with industrial demands, the practical application of these new technologies progresses relatively slowly. For example, one practitioner stated, “Industry is faster than academia at some things that lead to higher technology readiness levels. Industry tends not to be good at starting low trl development. In my experience, academia often develops things with little to no industrial application.” This highlights the advantage of industry in applying technology, especially in fields with higher Technology Readiness Levels (TRLs), where industry’s pace of advancement is faster. However, academia often conducts research at low TRL stages, leading to a disconnect between certain academic studies and practical industrial applications.
Additionally, 38.5% of practitioners believe that there are differences between academia and industry in the integration methods of LLMs and UAVs. 67.3% of practitioners think that academia and industry focus on different directions. Some practitioners pointed out: “Research - one topic, but deep, with logical knowledge base and understanding, industry - a lot of topics skimmed, practical usage with ‘black-box-magic”’. These statements illustrate that academia typically focuses on deeply solving specific problems and exploring unknowns and theoretical innovations, while industry is more concerned with practical applications, focusing on how to quickly, efficiently, and reliably apply technology.
Table 3. Practitioners’ perceptions on the reasons for differences
Challenge Definition Proportion
Technical Actual projects require mature and stable technology (75.0%) 84.6%
-Maturity Difficulty in actual project development and integration (55.8%)
Immaturity of lightweight large models (40.4%)
Performance High computing resource requirements (32.7%) 48.1%
Conflict between latency and real-time requirements (34.6%)
Safety Model robustness and reliability challenges (30.7%) 42.3%
Data privacy and security issues (19.2%)
Cost High computing power/hardware costs (46.2%) 44.2%
Commercial input-output ratio factors (28.8%)
Other Academic research is out of touch with industrial needs (36.5%) 57.7%
Regulatory and policy uncertainty (21.2%)
Lack of large model training data (25.0%)
Solutions to Bridge the Academia-Industry Gap. Furthermore, we asked practitioners to identify the reasons behind these differences. The responses can be categorized into the five challenges listed in  3. Among them, 84.6% of participants identified “Technical Maturity” as the main challenge. This indicates that, unlike academic research, which often emphasizes theoretical innovation, industry places greater importance on mature and reliable technical solutions. Specifically, practitioners pointed to the need to address issues such as “Actual projects require mature and stable technology” (75.0%), “Difficulty in actual project development and integration” (55.8%), and “Immaturity of lightweight large models” (40.4%). These findings reflect the challenges academic research faces in translating results into real-world applications, as industry often encounters complex conditions during technology integration and project development.
Practitioners also noted that industry faces challenges related to “Performance” (48.1%), “Safety” (42.3%), “Cost” (44.2%), and “Other” (57.7%). Corresponding issues include “High computational resource requirements” (46.2%), “Conflicts between latency and real-time requirements” (34.6%), “High computation/hardware costs” (32.7%), "Model robustness and reliability challenges" (30.8%), and “Commercial return on investment” (28.8%). One practitioner emphasized: “Industry wants practical results fast, especially smaller startups, etc. Academia has smaller groups to work with, less money, but great research skills and expert knowledge (even if not always practical).” This highlights the urgent demand for rapid results in industry, particularly in the process of commercialization, which makes the transfer of academic research outcomes into practical industrial applications more challenging.
Finding 12: Practitioners believe that the differences between academia and industry primarily lie in their focus areas (67.3%) and technical requirements (63.5%). The greatest challenge in addressing these differences is technical maturity (84.6%). Based on this, we can infer that although academic research can drive technological innovation, the development and integration of technology in practical projects face higher complexity and challenges. This necessitates consideration of the maturity, stability, and practical feasibility of the technology. This suggests that more attention should be paid to the alignment between academia and industry in terms of technology transfer and real-world application.
[…]
6. Discussion
This section first explains the significance of this study for researchers and practitioners, then summarizes the three main challenges of UAV–LLM applications, and finally presents the validity threats.
6.1. Implications
For researchers. This study reveals the significant differences between academia and industry in the integration of UAVs with LLMs. Academia typically focuses on theoretical modeling and the exploration of new technologies (e.g., 52.7% using External-Offboard Invocation modes), while industry is more concerned with practical applications and system integration (e.g., 50% using Hybrid-Onboard Assistive modes). Researchers should pay more attention to the technological needs in real-world applications, especially regarding real-time performance, robustness, and safety. Researchers should focus on developing lightweight models that can operate with low latency and high efficiency (Li et al., 2025a), particularly for complex tasks such as Multi-UAV Task Reasoning and Planning and Control Logic and Command Generation, optimizing algorithms to improve the accuracy and performance. Additionally, industry developers are concerned with the security and interpretability of large models. Researchers should enhance studies on model transparency and interpretability, developing technologies that ensure data privacy and security in critical tasks, thereby better supporting the widespread application of UAVs.
For practitioners. This study clarifies the tasks that current LLM technologies can support more reliably, such as task planning, user interaction, and automatic report generation. These tasks show stronger performance and lower implementation risks. The comparative analysis also indicates that many complex academic solutions are not yet ready for real-world use because of issues with latency, cost, and reliability. The results further point out important challenges in real-time processing and computational efficiency, giving developers a more realistic view of short-term priorities. To this end, practitioners can move forward in two directions. First, they can apply LLMs in tasks that are already feasible, such as building more natural user interfaces, generating reports automatically, and supporting mission planning. These areas involve lower risks and can deliver practical value relatively quickly. Second, for tasks that are not yet suitable for deployment, developers can focus on researching and testing improvements. This includes integrating lighter models, developing smaller or domain-specific models, or combining LLMs with traditional UAV control systems. In this way, practitioners can benefit from the current capabilities of LLMs while also preparing step by step for more advanced applications in the future.
6.2. Challenges of LLM–UAV Applications
(1) Bridging the Academia-Industry Gap in UAV-LLM Integration: There is a mismatch between academic research and industrial solutions. According to our survey, 67.3% of practitioners believed that academia and industry focus on different directions, and 63.5% thought that their technical requirements differ. The main reason is that academic research often concentrates on theoretical problems and overlooks industrial needs for real-time performance, stability, and system integration (75%). Industry, on the other hand, focuses on commercialization and rapid deployment, which slows down the transfer of academic outcomes to practice. To accelerate technology transfer, future research should better align with industrial needs. Researchers can collaborate with industry to obtain practical data and feedback, ensuring that new technologies can be effectively deployed in real-world environments and avoiding unnecessary delays and obstacles during technology transfer.
(2) Improving safety to increase practitioner trust in LLMs: Deploying LLMs on UAVs faces obstacles related to safety and reliability. Our survey shows that 54.8% of teams believe existing solutions are sufficient, while 45.2% highlight the security risks of large models. One practitioner explicitly stated: “LLMs should not be on a critical system due to their ‘black box’ nature. They are largely unneeded also.” These concerns stem from the lack of interpretability and transparency in LLMs, which can lead to uncertainty in decision-making and increase the risk of system failures. To improve the suitability of LLMs for UAV systems, future research should focus on enhancing interpretability and transparency, developing models that can explain their decision processes. In addition, developers should be provided with effective tools to monitor, validate, and debug model behavior, ensuring alignment with real task requirements, especially in complex and high-risk scenarios.
(3) Enhancing LLM Capabilities for Complex and Robust Autonomous Operations: Current LLMs are not sufficient to handle high-risk and long-duration autonomous UAV missions. As observed in Finding 8, industry interest in using LLMs for direct control execution is low (score: 2.05), with a strong preference for supportive roles (82.7%). Some practitioners further emphasized that “AI should never be put in direct control of a uav.” This reflects a reliance on traditional control algorithms rather than LLMs for decision-making, since current LLMs are viewed as unable to meet performance requirements (51.6%), and their capability in control logic generation remains immature (lowest score: 2.51). Although large-scale LLMs offer advantages in many complex tasks, their high computational resource demands limit their use in UAV systems, especially those requiring real-time performance with hardware constraints. One practitioner noted: “Lightweight large model research still lacks clear quantitative benchmarks for key trade-offs, so careful consideration is needed in model selection and distillation.” Therefore, future work should focus on developing efficient, lightweight models with lower computational demands to meet the real-time processing needs of UAV platforms. In addition, lightweight models should also be designed to handle the computational requirements of complex tasks such as multi-UAV coordination and task planning, ensuring stable and efficient operation in large-scale missions.
6.3. Threats To Validity
6.3.1. Internal Validity
During the process of literature screening and classification, researchers’ subjective judgments may influence the inclusion criteria and labeling of studies, potentially introducing bias into the analysis. To mitigate this threat, we implemented a dual-reviewer protocol across all stages of the analysis. Two authors independently conducted literature selection and labeling, and any disagreements were thoroughly discussed and resolved through consensus. It is worth noting that both authors involved in this process have more than two years of research experience in the field of UAVs, which helps ensure the validity and consistency of the analysis. In the survey component, practitioners may have provided dishonest or unprofessional answers. To reduce this risk, we explicitly informed participants at the beginning of the questionnaire that no personally identifiable information would be collected and that the survey was conducted anonymously. Except for the first four questions, all others were multiple-choice items, with the option to select “I don’t understand this question” for unclear items. Practitioners were free to skip any questions they found ambiguous. This design aimed to improve data quality while protecting practitioners’ privacy. According to Ong and Weiss (2000) (Ong and Weiss, 2000), ensuring anonymity and confidentiality in surveys is effective in eliciting unbiased responses. To collect more responses and overcome language barriers, we translated the questionnaire into Chinese and distributed it via platforms accessible in mainland China. While we made every effort to ensure consistency between the English and Chinese versions, minor discrepancies may have arisen during translation, which could affect practitioners’ understanding. Moreover, as all responses were written in Chinese, translation into English was required during data analysis, potentially introducing interpretation errors. To minimize this risk, two Chinese authors with strong English proficiency collaboratively reviewed and repeatedly refined the translations to ensure their accuracy and consistency.
6.3.2. External Validity
[…]
7. Related Work
7.1. UAV-LLM Integration
A growing body of research has investigated concrete systems and technical frameworks for integrating LLMs with UAVs. Chen et al.(Chen et al., 2023) developed TypeFly, the first real-world UAV platform integrated with an LLM to support natural language-driven flight control. Their system demonstrated the feasibility of mapping natural language to motion commands but also revealed limitations in handling semantic ambiguity and adapting to dynamic environments. Phadke et al.(Phadke et al., 2024) proposed a modular UAV control framework based on LLMs in simulation, coordinating navigation, obstacle avoidance, and target recognition modules. While the modular design improved stability, the absence of a unified semantic memory layer restricted decision efficiency. Zhao and Lin (Zhao et al., 2023b) introduced the concept of a Universal Aerial Agent with the “Prompt–as–Policy” approach, enabling flexible encoding of task intents, though their work remains at a conceptual stage without large-scale validation. More recently, Koubaa (Koubaa et al., 2025) introduced UAV–GPT, which translates high-level semantic instructions into executable UAV tasks, while Ren et al. (Ren et al., 2024) integrated foundation models with UAV perception for semantic annotation and visual reasoning.
These technical explorations highlight the potential of LLMs to enhance UAV autonomy, perception, and interaction, yet most remain isolated prototypes or proof-of-concept systems. Compared to these works, our study provides a more comprehensive and integrative perspective: we categorize task roles across the UAV workflow, empirically compare academic versus industrial tendencies, and identify bottlenecks from practitioner data. This broader view helps to contextualize existing prototypes within a systematic landscape and points out where research and practice can be better aligned.
7.2. Surveys on UAVs and Machine Learning
A number of surveys have summarized the use of machine learning techniques in UAVs. For example, Osco et al.(Osco et al., 2021) reviewed the role of deep learning in UAV remote sensing, covering tasks such as image classification, object recognition, and environmental monitoring. Tang et al.(Tang et al., 2023) provided a survey on object detection for UAVs based on deep learning, emphasizing algorithmic advances and benchmark datasets. AlMahamid and Grolinger (AlMahamid and Grolinger, 2022) systematically reviewed reinforcement learning approaches for autonomous UAV navigation, highlighting progress and challenges in dynamic environments. Bai et al. (Bai et al., 2023) focused on reinforcement learning–based methods for building autonomous multi-UAV wireless networks, summarizing techniques for coordination and communication. These surveys clearly show the important role of machine learning techniques in UAV research, but none of them address the emerging opportunities and challenges brought by large language models. In other words, although UAV–DL/ML integration has been extensively reviewed, the UAV–LLM integration landscape remains largely unexplored.
There has been one recent attempt to discuss UAV-LLM integration. Javaid et al. (Javaid et al., 2024) reviewed possible applications of LLMs in UAVs, focusing on natural language command control, reasoning-assisted decision making, and human-computer interaction. However, their work mainly focuses on conceptual and descriptive descriptions, without connecting academic, industry, and developers’ perspectives. In contrast, our study constructs the first taxonomy of UAV-LLM tasks, analyzes 997 academic papers and 1,509 open-source projects, and further validates our findings with survey responses from 52 developers. This large-scale effort not only summarizes existing research but also quantifies the gap between academia and industry, offering practical guidance for real-world deployment.
[…]
8. Conclusion
In this paper, we conduct an empirical study on the integration of LLMs and UAVs to examine the current state of applications in academia and industry. First, by analyzing 74 papers and 56 open-source projects on GitHub, we summarize nine categories of LLM-supported tasks across four UAV workflows. Second, we quantify the distribution of these tasks in academic and industrial settings and perform a comparative analysis. The results reveal gaps between academia and industry in research focus, technical requirements, and integration modes of LLMs and UAVs. We then administer an online survey and obtain 52 valid responses. We find that 59.6% of practitioners have not yet attempted to use LLMs in real-world UAV projects. Their feedback indicates that current LLMs fall short of meeting the requirements of UAVs for accuracy, real-time performance, and reliability, and that the lack of technological maturity and practical feasibility in academic research further increases the complexity of engineering deployment. Finally, based on our findings, we highlight future challenges for UAV-LLM integration and offer recommendations for researchers and practitioners.